OpenStudy (anonymous):

In order to make 100 mL of a 0.512 M NaCl solution, then _________ g of NaCl is needed.

OpenStudy (taramgrant0543664):

The formula is m=cv Where m is the mass that you are determining, c is the concentration in this case 0.512M and v is volume (note volume is in litres so you would have to convert it)

OpenStudy (anonymous):

so is it 2.99g ?

OpenStudy (taramgrant0543664):

100mL=0.1L m=cv =(.512)(.1) =0.0512g

OpenStudy (anonymous):

but we're trying to make a 100mL solution of 0.512g how is the answer the same as the number in the question with one more 0???

OpenStudy (anonymous):

0.512M*

OpenStudy (taramgrant0543664):

Because it's 100mL solution (or 0.1L) it would change more if it was at a different volume like if it was at 200mL you would have 0.102g of NaCl in it

OpenStudy (anonymous):

So is the answer 2.99?

OpenStudy (anonymous):

What I did was (0.1L x 58.4 x 0.512) because there is 58.4 g/mol in NaCl

OpenStudy (taramgrant0543664):

You don't need to put the molar mass in at all Let's say if you did put in 2.99g of NaCl c=m/v =2.99/0.1 =29.9 as your concentration